# Hadamard Attention Recurrent Transformer: A Strong Baseline for Stereo Matching Transformer

### v2

@@ III-D Simple Gate Feed Forward (SGFF)
where $GELU$ means GELU operation [34]. The overall computation process of our transformer block can be represented as Equ. 9.
@@
where $Block$ means the our transformer block in Fig. 2, $LN$ denotes to Layer Normalization [35]. $Q$, $K$, and $V$ here are three variables derived from the initial feature $V^{init}$. A 1$\times$1 convolutional layer is employed to produce features $V^{3c}$ with a channel size of $3c$. These features are then evenly split into three parts along the channel dimension, representing $Q$, $K$, and $V$ respectively.
III-E IGEV Decoder
@@
[…]
TABLE I: Default hyperparameters of HART.
@@
Figure 5: Visualisation on the Middlebury dataset [39]. We conducted comparisons with existing SOTA methods [8, 40]. Disparity and error maps at 0.5 px threshold are included.
TABLE II: Quantitative evaluation on Scene Flow test set, the 1-pixel error rate is employed (lower is better). The best results are indicated in bold and the second-best results are underscored. $\star$ refers to Stereo Transformer methods.
Method Publish EPE (px)$\downarrow$ D1 (%$\downarrow$
GC-Net [41] ICCV2017 1.84 15.6
PSMNet [14] CVPR2018 1.09 12.1
GANet [42] CVPR2019 0.78 8.7
GwcNet [15] CVPR2019 0.76 8.0
AANet [43] CVPR2020 0.87 9.3
RAFT-Stereo [6] 3DV2021 0.61 6.5
ACVNet+ [32] CVPR2022 0.48 5.0
IGEV-Stereo [8] CVPR2023 0.47 5.3
UPFNet [44] TCSVT2023 0.71 6.7
DMCA-Net [45] TCSVT2023 0.64 6.9
Selective-IGEV [46] CVPR2024 0.44 5.0
STTR [11] $\star$ ICCV2021 0.45 -
CroCo-Stereo [21] $\star$ ICCV2023 - 5.3
DLNR [22] $\star$ CVPR2023 0.48 5.4
GOAT [47] $\star$ WACV2024 0.47 5.6
HART (Ours) $\star$ Ours 0.42 5.0
TABLE III: Results on the KITTI 2012 Reflective leaderboard. “Noc” means percentage of erroneous pixels in non-occluded areas, “All” means percentage of erroneous pixels in total. Bold: best performance, underline: second best.
@@
TABLE IV: Results on the KITTI 2015 and 2012. Error threshold is 3 px for 2015, and 4 px for 2012. Front-ground error is indicated by ”F.G.”, and Background error by ”B.G.”. $\star$ refers to Stereo Transformer methods, which are summarized in the lower half of the table. The upper half of the table describes Stereo CNN methods. “Time” denotes the inference time on single NVIDIA Tesla A100.
@@
Selective-IGEV [46] CVPR2024 2.61 1.33 2.55 1.22 0.84 1.07 0.24
STTR [11] $\star$ CVPR2021 3.61 1.70 - - - - 0.88
@@
IV Experiment
IV-A Implementation Details
The PyTorch framework is used to implement HART. The AdamW optimizer [56] is used during training. Tab. I collates the hyperparameters used by HART. The KITTI 2012 [57], KITTI 2015 [31], Scene Flow [58], and Middlebury [39] datasets were used to assess the performance of our model. HART is trained for $200,000$ steps on the Scene Flow [58] dataset with a batch size of $8$. The input photographs are cropped at random to 320 $\times$ 736 pixels. We use data augmentation methods including spatial transformations and asymmetric chromatic augmentations. For training on the Middlebury dataset [39], we begin by fine-tuning the Scene Flow [58] pretrained model with a crop size of 384 $\times$ 512 for 200,000 steps, using the mixed Tartan Air [59], CREStereo [16], Scene Flow [58], Falling Things [60], InStereo2k [61], CARLA [62], and Middlebury [39] datasets. We then use a batch size of 8 for an additional 100,000 steps to refine the model on the mixed CREStereo dataset [16], Falling Things [60], InStereo2k [61], CARLA [62], and Middlebury datasets [39]. For training on the KITTI 2012 [57] and KITTI 2015 [31] datasets, we fine-tune the Scene Flow pretrained model on the combined KITTI 2012 and KITTI 2015 datasets for 50,000 steps.
IV-B Comparisons with State-of-the-art
@@
Scene Flow. As shown in Tab. II, HART achieves SOTA performance on the Scene Flow dataset. It achieves a 10.6% error reduction on the EPE metric compared to GOAT [47], the SOTA Stereo Transformer, and by 6.0% on the D1 metric compared to CroCo-Stereo [21].
Reflective Area. Due to the stronger capability in capturing long-range information, HART can integrate global features to make more accurate matches in ill-posed regions, such as reflective area. HART ranked 1st on the KITTI 2012 reflective benchmark at the time of submission, as shown in Tab. III.
KITTI. To evaluate the performance of HART in real-world scenarios, we conducted experiments on KITTI benchmarks. Stereo transformers exhibit a significant disadvantage in terms of time. This is primarily due to the quadratic time complexity of their SA. Although DLNR significantly outperforms other stereo transformers in terms of inference speed, it sacrifices spatial dimension interaction. Experimental results in Tab. IV and Fig. 4 confirm that, HART is still able to emphasise the matching of spatial details within a given inference time. While compared to SOTA methods, HART does exhibit some deficiencies in background matching. This attributed to our designs, which leads to focus more on foreground features.
TABLE V: Results on Middlebury. The error threshold are 0.5 px and 1 px here.
@@
Figure 6: Zero-shot evaluation without fine-tuning on Middlebury. All results visualised here are trained on Scene Flow. The odd-numbered columns show the original images, the even-numbered columns present zoomed-in details for better visualization.
TABLE VI: Zero-shot evaluation on Middlebury and KITTI. Every model undergoes scene flow training without fine-tuning. The 2px error rate is employed for Middlebury, and 3px error rate for KITTI.
@@
HART Ours 12.2 4.9 4.8 1.1 5.1
TABLE VII: Zero-shot evaluation on the Driving Stereo dataset. The metric used is the End Point Error (EPE) on the full-resolution (F) and half-resolution (H) images.
@@ F H F H F H F H
TABLE VIII: Ablation study. Error threshold is 3 px. “ID” means IGEV Decoder, “HP” denotes to the Hadamard Product between Query and Key here. The following metrics represent the results obtained from training and testing on the Scene Flow dataset. “Time” denotes the inference time on single NVIDIA Tesla A100.
No. STTR HPSA Encoder ID EPE D1 Time
Encoder HP SoftMax DAK MKOI (px) (%) (s)
1 - - - - - - 0.47 2.47 0.37
2 0.44 2.38 0.71
3 0.49 2.46 0.28
4 0.46 2.43 0.28
5 0.43 2.39 0.36
6 0.42 2.32 0.36
@@
Middlebury. HART is also able to achieve excellent performance in the Middlebury scenario. HART outperforms the SOTA method LoS [40] by 15.5%, 15.1%, 8.62% and 11.3% for the four metrics shown in Tab. V and Fig. 5.
IV-C Zero-shot Generalization
@@
Collecting stereo matching datasets poses significant challenges, and stereo matching in real-world scenes often requires training on a small dataset to obtain matching results in untrained real scenes. This places certain demands on the generalization performance of stereo matching models. Zero-shot generalization is a common experiment to test the generalization performance of stereo matching. We follow the paradigm of training on Scene Flow and testing the model on real scenes from Middlebury, KITTI, and Driving Stereo [65] datasets to validate our generalization ability. The experimental results are presented in Tab. VI, Tab. VII and Fig. 6. The quantitative comparison results demonstrate the excellent generalization performance of HART.
IV-D Ablations
@@
To validate the effectiveness of our design, we conducte ablation studies by individually removing each module. All experiments are conducted on the Scene Flow [58] dataset, and all hyperparameters are kept consistent with [8]. The results of the ablation experiments are shown in Tab. VIII.
@@
Baseline No.1: Recurrent Stereo CNN. Our No.1 baseline is IGEV-Stereo [8]. Experiment No.1 in Tab. VIII demonstrates the performance of this method. Our transformer decoder borrows its code related to iterative updating and correlation calculation.
@@
Baseline No.2: Recurrent Stereo Transformer. Building upon STTR [11], we designed the STTR encoder with the IGEV Decoder, as introduced in Section III-E. The improved version of STTR (No. 2 Experiment) has achieved higher accuracy and faster inference time compared to the original STTR [11]. This improvement is attributed not only to the enhanced decoder, i.e., IGEV Decoder, but also to the adoption of techniques inspired by Restormer [23]. Specifically, the feed-forward propagation network of the encoder has been upgraded to the SGFF structure introduced in Section III-D. Through this combination, we obtained a Recurrent Stereo Transformer. The effectiveness of this Recurrent Stereo Transformer, which is based on the vanilla self-attention (SA), is showcased in Experiment No.2.
@@
TABLE IX: Ablation study for iterations. The metric here is end point error
@@
HART(Ours) 0.60 0.55 0.51 0.48 0.45 0.42
TABLE X: Zero-shot performance of HART-DLNR. The baseline here is DLNR. Our proposed HPSA Is used to replace self-attention mechanism (CATE) of DLNR. All models are trained on Scene Flow, and tested on Middlebury and KITTI. The threshold is 2px for Middlebury, and 3px for KITTI. “Time (s)” is the inference time required for the model to run on the KITTI dataset using the NVIDIA A100.
@@
Hadamard Product between Query and Key. Hadamard product (HP) is the foundation of our HPSA. By utilizing the Hadamard product for attention computation, HART achieves linear attention computation. The effectiveness of this caculation is demonstrated in Experiment No.3. Although combining HP with the vanilla SoftMax activation function significantly reduces inference time, this approach encounters two issues. 1) The limitation of the SoftMax activation function prevents HART from overcoming the low-rank bottleneck [13] problem. 2) The Hadamard product lacks spatial and channel interactions, making it challenging to achieve satisfactory performance solely through the use of the Hadamard Product.
Dense Attention Kernel (DAK). DAK is proposed to address the low-rank bottleneck issue caused by SoftMax. Assigning larger differential activation values to the attention matrix can also make the matching features more precise. Comparative experiments No.3 and No.4 demonstrate the effectiveness of DAK.
$HPSA(Q,K,V)_{w/o~{}MKOI}=DAK(||Q||_{2}\odot||K||_{2})\odot V$ (11)
It is difficult to achieve spatial and channel feature interaction with this method, resulting in information loss. We developed MKOI to overcome this shortcoming. Experiments No.5 and No.6 demonstrate the performance of MKOI.
Number of Iterations We adopted standard parameter settings and conducted our tests with 32 iterations. The architecture of HART enhances the effectiveness of the iterative strategy. HART achieves superior performance compared to other recurrent-based stereo methods across various iteration counts. To substantiate this claim, we compared HART against the SOTA Recurrent Stereo Transformer [22] and Recurrent Stereo CNN [8]. As shown in Tab. IX, experiments demonstrate that HART consistently outperforms other recurrent-based methods across various iteration counts.
@@
IV-E Extension to existing Stereo Transformer
To further validate that our attention paradigm is a better way for stereo transformers, we applied HPSA from our HART to replace the Channel-Attention Transformer Extractor (CATE) from DLNR. DLNR is a representative Recurrent Stereo Transformer. Although CATE makes less inference time than other stereo transformers [12, 21], the trade-off is the loss of classical spatial interactions achieved by traditional ViTs. What is more, it still suffers from common issues inherent to stereo transformers. Firstly, CATE still exhibits quadratic time complexity, leaving room for improvement in inference time. Additionally, the parameter count for each attention head in CATE is still $2nc/d$, which remains significantly smaller than $c^{2}$. This implies that DLNR also faces the issue of low-rank bottleneck [13]. Compared to CATE in DLNR, HPSA offers the following advantages: 1) Further reduction in computational complexity; 2) Alleviating the low-rank bottleneck by using DAK; 3) Spatial dimension interaction has been achieved; 4) Improve the way channel dimensions interact. These benefits address the limitations of CATE and allow HPSA to replace CATE and achieve better outcomes. Experimental results shown in Fig. 7 and Tab. X demonstrate that HPSA designs improve the performance of DLNR and reduce the inference time.
V Conclusion
HART represents a novel approach to the transformer architecture. The objective is to surmount the intrinsic constraints of the square-level complexity and low-rank bottleneck, which have an adverse impact on the efficacy of stereo matching transformers. HART controls the time complexity within O(n) using the Hadamard product. DAK is established to focus on important features and to ensure the potential for rank growth. MKOI is proposed to complement the spatial and channel interaction capabilities. These designs allow HART to achieve SOTA performance. Experimental results demonstrate the effectiveness of our HART. HART ranks 1st on the KITTI 2012 reflective benchmark and TOP 10 on the Middlebury benchmark among all published methods. While HART still has its limitation. HART is not able to achieve real-time at the moment. In the future, we plan to implement a faster solution and validate the effectiveness of HART in more frameworks.
Acknowledgment

### v4

@@ 3.4 Simple Gate Feed Forward (SGFF)
where $GELU$ means GELU operation. The overall computation process of our transformer block can be represented as Equ. 9.
@@
where $Block$ means the our transformer block in Fig. 2, $LN$ denotes to Layer Normalization [35]. $Q$, $K$, and $V$ here are three variables derived from the initial feature $V^{init}$. A $\times1$ convolutional layer is employed to produce features $V^{3c}$ with a channel size of $3c$. These features are then evenly split into three parts along the channel dimension, representing $Q$, $K$, and $V$ respectively.
3.5 IGEV Decoder
@@
where $N_{c}$ is the number of channels, $N_{g}$ means the group number ($N_{g}=8$ here), $d$ denotes to the disparity level, $\langle\cdots,\cdots\rangle$ is the inner product, $F$ is a frozen network pretrained by [38], $l$ is the left view image. The construction of 3D regularization network $\mathbb{R}$ also follows IGEV-Stereo [7].
Table 1: Default hyperparameters of HART.
@@
Figure 5: Visualisation on the Middlebury dataset [39]. Disparity and error maps at 0.5 px threshold are included. In the error maps, black regions indicate mismatched areas. We conducted comparisons with existing SOTA methods [7, 40]. It can be observed that HART produces fewer mismatched regions compared to other methods.
Table 2: Quantitative evaluation on Scene Flow test set, the 1-pixel error rate is employed (lower is better). The best results are indicated in bold. $\star$ refers to Stereo Transformer methods.
[…]
GC-Net [41] ICCV2017 1.84
PSMNet [13] CVPR2018 1.09
GANet [42] CVPR2019 0.78
GwcNet [14] CVPR2019 0.76
AANet [43] CVPR2020 0.87
RAFT-Stereo [5] 3DV2021 0.61
ACVNet+ [33] CVPR2022 0.48
IGEV-Stereo [7] CVPR2023 0.47
Selective-IGEV [44] CVPR2024 0.44
STTR [10] $\star$ ICCV2021 0.45
DLNR [20] $\star$ CVPR2023 0.53
GOAT [45] $\star$ WACV2024 0.47
RetinaStereo [21] $\star$ ICASSP2025 0.49
HART (Ours) $\star$ Ours 0.42
Table 3: Results on the KITTI 2012 Reflective leaderboard. “Noc" means percentage of erroneous pixels in non-occluded areas, “All" means percentage of erroneous pixels in total.
@@
Table 4: Results on the KITTI 2015 and 2012. Error threshold is 3 px for 2015, and 4 px for 2012. Front-ground error is indicated by "F.G.", and Background error by "B.G.". $\star$ refers to Stereo Transformer methods, which are summarized in the lower half of the table. The upper half of the table describes Stereo CNN methods. “Time" denotes the inference time on single NVIDIA Tesla A100.
@@
STTR [10] $\star$ CVPR2021 3.61 1.70 - - - - 0.78
@@
[…]
4 Experiment
4.1 Implementation Details
The PyTorch framework is used to implement HART. The AdamW optimizer is used during training. Tab. 1 collates the hyperparameters used by HART. The KITTI 2012 [52], KITTI 2015 [32], Scene Flow [53], and Middlebury [39] datasets were used to assess the performance of our model. HART is trained for $200,000$ steps on the Scene Flow [53] dataset with a batch size of $8$. The input photographs are cropped at random to 320 $\times$ 736 pixels. We use data augmentation methods including spatial transformations and asymmetric chromatic augmentations. For training on the Middlebury dataset [39], we begin by fine-tuning the Scene Flow [53] pretrained model with a crop size of 384 $\times$ 512 for 200,000 steps, using the mixed Tartan Air [54], CREStereo [15], Scene Flow [53], Falling Things [55], InStereo2k [56], CARLA [57], and Middlebury [39] datasets. We then use a batch size of 8 for an additional 100,000 steps to refine the model on the mixed CREStereo dataset [15], Falling Things [55], InStereo2k [56], CARLA [57], and Middlebury datasets [39]. For training on the KITTI 2012 [52] and KITTI 2015 [32] datasets, we fine-tune the Scene Flow pretrained model on the combined KITTI 2012 and KITTI 2015 datasets for 50,000 steps.
4.2 Comparisons with State-of-the-art
@@
Scene Flow. As shown in Tab. 2, HART achieves SOTA performance on the Scene Flow dataset. It achieves a 6.67% error reduction on the EPE metric compared to STTR [10].
Reflective Area. HART can integrate features to make more accurate matches in ill-posed regions, such as reflective area. HART ranked 1st on the KITTI 2012 reflective benchmark at the time of submission, as shown in Tab. 3.
KITTI. To evaluate the performance of HART in real-world scenarios, we conducted experiments on KITTI benchmarks. Stereo transformers exhibit a significant disadvantage in terms of time. This is primarily due to the quadratic time complexity of their self-attention. While DLNR achieves superior inference speed compared to other stereo transformers, this efficiency comes at the cost of compromised spatial dimension interaction. Consequently, the model exhibits constrained stereo matching accuracy in dynamic driving environments. HART overcomes this limitation and achieves improved performance under constrained inference time. Experimental results in Tab. 4 and Fig. 4 confirm that, HART is still able to emphasise the matching of spatial details within a given time.
Table 5: Results on Middlebury. The error threshold are 0.5 px and 1 px here.
@@
Figure 6: Zero-shot evaluation without fine-tuning on Middlebury. All results visualised here are trained on Scene Flow. The odd-numbered columns show the original images, the even-numbered columns present zoomed-in details for better visualization. In the first scenario, IGEV-Stereo misclassifies industrial equipment components as background and fails to recognize the thermometer as a co-planar object. In the second scenario, IGEV-Stereo is confused by thin objects, leading to incorrect disparity estimation in regions such as gaps within baskets and chairs.
Table 6: Zero-shot evaluation on Middlebury and KITTI. Every model undergoes scene flow training without fine-tuning. The 2px error rate is employed for Middlebury, and 3px error rate for KITTI.
@@
STTR [10] $\star$ ICCV2021 - 15.5 9.7 1.5 6.4
CSTR [58] $\star$ ECCV2022 - - - 1.4 5.8
ELFNet [11] $\star$ ICCV2023 - 10.0 7.9 1.6 5.8
DLNR [20] $\star$ CVPR2023 14.5 9.5 7.6 2.8 16.1
FET [51] $\star$ TCI2024 - - - 1.4 6.7
HART $\star$ Ours 12.2 4.9 4.8 1.1 5.1
Table 7: Zero-shot evaluation on the Driving Stereo dataset. The metric used is the End Point Error (EPE) on the full-resolution (F) and half-resolution (H) images.
@@ F H F H F H F H
Table 8: Ablation study. Error threshold is 3 px. “ID" means IGEV Decoder. The following metrics represent the results obtained from training and testing on the Scene Flow dataset. $\frac{Rank(A)}{m}$ computes the average percentage of the ratio between the rank of the attention matrix and its dimension $m$ (the number of rows in the attention matrix) across inference steps. “Time" denotes the inference time on single NVIDIA Tesla A100.
No. STTR HPSA Encoder ID $\frac{R(A)}{m}$ EPE D1 Time
Encoder Attention MKOI (%) (px) (%) (s)
1 CNN architecture 0.47 2.47 0.37
2 Vanilla SA - 78.1 0.44 2.38 0.71
3 - Hadamard product w/ SoftMax - 81.3 0.49 2.46 0.28
4 - Hadamard product w/ DAK - 93.0 0.46 2.43 0.28
5 - Hadamard product w/ SoftMax 81.8 0.43 2.39 0.36
6 - Hadamard product w/ DAK 94.5 0.42 2.32 0.36
Table 9: Ablation studies on the MKOI module. Experiment No. 1 corresponds to the default configuration of MKOI. The table under "Value Matrix" summarizes the convolution operations along with their corresponding input and output channel dimensions.
No. Attention Matrix Value Matrix EPE D1 Para. (M)
$\times3$ $\times5$ $\times7$
1 $\times1$ Conv, $\frac{7}{4}$C C C/2 C/4 0.42 2.32 107.27
2 C C C 0.43 2.31 163.13
3 C/4 C/2 C 0.45 2.48 151.01
4 $\times7$ Conv, $\frac{7}{4}$C C C/2 C/4 0.43 2.40 170.68
5 $\times1$ Conv, $\frac{7}{4}$C Only $\times3$ Conv, $\frac{7}{4}$C Only 0.46 2.44 96.12
@@
Middlebury. HART is also able to achieve excellent performance in the Middlebury scenario. HART outperforms the SOTA method LoS [40] by 15.5%, 15.1%, 8.62% and 11.3% for the four metrics shown in Tab. 5 and Fig. 5.
4.3 Zero-shot Generalization
@@
Collecting stereo matching datasets poses significant challenges, and stereo matching in real-world scenes often requires training on a small dataset to obtain matching results in untrained real scenes. This places certain demands on the generalization performance of stereo matching models. Zero-shot generalization is a common experiment to test the generalization performance of stereo matching. We follow the paradigm of training on Scene Flow and testing the model on real scenes from Middlebury, KITTI, and Driving Stereo [60] datasets to validate our generalization ability. The experimental results are presented in Tab. 6, Tab. 7 and Fig. 6. The quantitative comparison results demonstrate the excellent generalization performance of HART.
4.4 Ablations
@@
To validate the effectiveness of our design, we conducte ablation studies by individually removing each module. All experiments are conducted on the Scene Flow [53] dataset, and all hyperparameters are kept consistent with [7]. The results of the ablation experiments are shown in Tab. 8. In addition to standard metrics such as End Point Error (EPE) and D1, we also evaluate the $\frac{Rank(A)}{m}$ metric. This new metric computes the average percentage of the ratio between the rank of the attention matrix and its dimension $m$ (i.e., the number of rows in the attention matrix) across inference steps.
@@
Baseline No.1: Recurrent Stereo CNN. Our No.1 baseline is IGEV-Stereo [7]. Experiment No.1 in Tab. 8 demonstrates the performance of this method. Our transformer decoder borrows its code related to iterative updating and correlation calculation.
@@
Baseline No.2: Recurrent Stereo Transformer. Building upon STTR [10], we designed the STTR encoder with the IGEV Decoder, as introduced in Section 3.5. The improved version of STTR (No. 2 Experiment) has achieved higher accuracy and faster inference time compared to the original STTR [10]. This improvement is attributed not only to the enhanced decoder, i.e., IGEV Decoder, but also to the adoption of techniques inspired by Restormer [22]. Specifically, the feed-forward propagation network of the encoder has been upgraded to the SGFF structure introduced in Section 3.4. Through this combination, we obtained a Recurrent Stereo Transformer. The effectiveness of this Recurrent Stereo Transformer, which is based on the vanilla self-attention (SA), is showcased in Experiment No.2.
@@
Table 10: Ablation study for iterations. The metric here is end point error
@@
DLNR [20] $\star$ 1.56 0.96 0.80 0.71 0.58 0.53
RetinaStereo [21] $\star$ 1.00 0.80 0.69 0.62 0.52 0.49
HART (Ours) $\star$ 0.60 0.55 0.51 0.48 0.45 0.42
Hadamard product between Query and Key. Hadamard product is the foundation of our HPSA. By utilizing the Hadamard product for attention computation, HART achieves linear attention computation. The effectiveness of this caculation is demonstrated in Experiment No.3. Although combining HP with the vanilla SoftMax activation function significantly reduces inference time, this approach encounters two issues. 1) The limitation of the SoftMax activation function prevents HART from overcoming the low-rank bottleneck [12] problem. 2) The Hadamard product lacks spatial and channel interactions, making it challenging to achieve satisfactory performance solely through the use of the Hadamard product.
Dense Attention Kernel (DAK). The SoftMax operation converts attention weights into a probability distribution through exponential normalization. However, this process excessively suppresses features at certain positions, driving some weights in the attention matrix close to zero and consequently leading to loss of features. As shown in Fig. 7 Such sparsity may impair the model’s ability to capture contextual information in ill-posed region, resulting in a low-rank bottleneck. DAK is proposed to address the low-rank bottleneck issue caused by SoftMax. The DAK ensures that the attention matrix remains positive, enabling effective focus on ill-posed regions. As shown in Tab. 8, from the perspective of matrix rank, DAK contributes to achieving higher ranks, which to some extent enhances the model’s representational capacity. Comparative experiments No.3 and No.4 demonstrate the effectiveness of DAK.
[…]
Since the interaction between matrices is performed entirely through Hadamard product, Equ. 11 fails to facilitate spatial and channel dimensional interactions. We further developed MKOI to overcome this shortcoming. Experiments No.5 and No.6 demonstrate the performance of MKOI.
We argue that semantic information has limited correlation with fine-grained details, and thus full feature channels are not necessary to achieve comparable performance. Consequently, we apply large-kernel convolutions only on a subset of the channels. Experiments No. 1 and No.2 employ $\times1$ convolutions to expand the feature channels of the attention matrix from C to $\frac{7}{4}$C. As shown in Tab. 9, when the channel dimension for large-kernel convolution is reduced to C/4, the number of parameters decreases while maintaining equivalent accuracy.
Local detail features require sliding convolution with small kernels across all channels. Experiment No. 3 adopts a design logic opposite to that of MKOI: large-kernel convolutions are applied to all feature channels, while small-kernel convolutions are used on a subset of channels. Experiment No. 4 utilizes a $\times7$ convolution to upscale the feature channels of the attention matrix from C to $\frac{7}{4}$C. These results demonstrate that using only a subset of channels or applying large-kernel convolutions on the attention matrix fails to achieve satisfactory performance.
Furthermore, experiment No. 5 employs only standard $\times1$ convolutions: first, to expand the feature channels of both the attention and value matrices from C to $\frac{7}{4}$C; then, their Hadamard product is passed through another $\times1$ convolution to reduce the channel dimension back to C, producing the final output of the attention process. The comparison with standard convolution further demonstrates the effectiveness of MKOI, and confirms that MKOI’s paradigm—splitting channels first and then performing interaction—is effective.
Number of Iterations We adopted standard parameter settings and conducted our tests with 32 iterations. The architecture of HART enhances the effectiveness of the iterative strategy. HART achieves superior performance compared to other recurrent-based stereo CNNs [5, 7, 44] and stereo transformers [20, 21] across various iteration counts. To substantiate this claim, we compared HART against the SOTA Recurrent Stereo Transformer [20] and Recurrent Stereo CNN [7]. As shown in Tab. 10, experiments demonstrate that HART consistently outperforms other recurrent-based methods across various iteration counts.
Figure 7: Visualization of the effects using DAK and SoftMax as kernel functions is presented. The figure shows a visualization of the model trained on the Scene Flow Train dataset, evaluated on a sample scene from the Scene Flow Test dataset. As highlighted in the zoomed-in region, using DAK as the kernel function leads to improved matching performance in ill-posed regions such as reflective surfaces, edge textures, and thin objects.
Table 11: Zero-shot performance of HART-DLNR. The baseline here is DLNR. Our proposed HPSA Is used to replace self-attention mechanism (CATE) of DLNR. All models are trained on Scene Flow, and tested on Middlebury and KITTI. The threshold is 2px for Middlebury, and 3px for KITTI. “Time (s)" is the inference time required for the model to run on the KITTI dataset using the NVIDIA A100.
@@
4.5 Extension to existing Stereo Transformer
To further validate that our attention paradigm is a better way for stereo transformers, we applied HPSA from our HART to replace the Channel-Attention Transformer Extractor (CATE) from DLNR. DLNR is a representative Recurrent Stereo Transformer. Although CATE makes less inference time than other stereo transformers [11, 17], the trade-off is the loss of classical spatial interactions achieved by traditional ViTs. What is more, it still suffers from common issues inherent to stereo transformers. Firstly, CATE still exhibits quadratic time complexity, leaving room for improvement in inference time. Additionally, the parameter count for each attention head in CATE is still $2nc/d$, which remains significantly smaller than $c^{2}$. This implies that DLNR also faces the issue of low-rank bottleneck [12]. Compared to CATE in DLNR, HPSA offers the following advantages: 1) Further reduction in computational complexity; 2) Alleviating the low-rank bottleneck by using DAK; 3) Spatial dimension interaction has been achieved; 4) Improve the way channel dimensions interact. These benefits address the limitations of CATE and allow HPSA to replace CATE and achieve better outcomes. Experimental results shown in Fig. 8 and Tab. 11 demonstrate that HPSA designs improve the performance of DLNR and reduce the inference time.
5 Conclusion
HART introduces a novel design of the transformer architecture. The goal is to improve the stereo matching capability of transformers in ill-posed regions while overcoming the limitations imposed by quadratic computational complexity. DAK is established to solve the intrinsic constraints of low-rank bottleneck. MKOI is proposed to complement the spatial and channel interaction capabilities. The overall pipeline controls the time complexity within $O(n)$ using an attention pipeline based on Hadamard product. These designs allow HART to achieve SOTA performance. Experimental results demonstrate the effectiveness of our HART. HART ranks 1st on the KITTI 2012 reflective benchmark and TOP 10 on the Middlebury benchmark among all published methods. While HART still has its limitation. HART is not able to achieve real-time at the moment. In the future, we plan to implement a faster solution and validate the effectiveness of HART in more frameworks.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Funding